OpenStudy (anonymous):

Simplify. Please show all of your work. 287-8/6-6^2

OpenStudy (anonymous):

Hollywood isn't written with a z and please use proper notation other wise your question is ambiguous

OpenStudy (anonymous):

287-8 = 279 6^(2) = 6*6 Remember BEDMAS 279/6-(6)(6) = 279/30

OpenStudy (anonymous):

you can probably simplify thus further by dividing the top and bottom by 3

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem, if you have anyother questions regarding algebra just ask